OpenStudy (anonymous):

Pre-cal A sector of a circle of radius 24 mi has an area of 288 mi^2. Find the central angle of the sector. Please HELP!

OpenStudy (saifoo.khan):

@OpenstudyUser

OpenStudy (anonymous):

I know the answer is 1 rad or 57.3 but idk how!

OpenStudy (saifoo.khan):

\[281 = \frac{x}{2 \pi} \times \pi \times 24^2\]

OpenStudy (openstudyuser):

@saifoo.khan I just started the course.

OpenStudy (saifoo.khan):

@OpenstudyUser , you're good for nothing. -_-

OpenStudy (openstudyuser):

Yea, i guess.

OpenStudy (anonymous):

i didnt get 57.3 though :/